# Learning Personalized Decision Support Policies

### v2

@@ Appendix B Comparison Against Prior Work
  • CONSENSUS: One can also depict forms of support that are independent of any model, for instance, presenting the belief of one or more humans. Belief distributions can be constructed by pooling over many different humans’ “votes” for what a label ought to be, e.g., (Peterson et al. 2019; Beyer et al. 2020; Uma, Almanea, and Poesio 2022; Uma et al. 2020; Gordon et al. 2021, 2022), or by eliciting distributions over the likely label directly from each individual human (Collins, Bhatt, and Weller 2022; Collins et al. 2023a). These consensus distributions permit the expression of uncertainty without any model. However, the elicitation of this form of support may be costly and humans may be fallible in the information they provide, e.g., due to direct labeling errors (Dawid and Skene 1979; Augustin et al. 2017; Whitehill et al. 2009; Wei et al. 2022), or miscalibrated confidence (O’Hagan et al. 2006; Collins et al. 2023a; Lichtenstein, Fischhoff, and Phillips 1977; Tversky and Kahneman 1996).
  • ADDITIONAL: While much of this paper focused on support that provides decision-makers with label information, decision support may also entail acquiring or displaying additional contextual information (e.g., new features (Bakker et al. 2021)) or requesting previously unseen features, for instance, through additional medical diagnostics  (Harrell et al. 1982; Mylonakis et al. 2000). This flavor of support can be varied structurally, ranging from the results of a search query (Nakano et al. 2021) to hierarchical information like exposing the subsidiary ownership structure for multinational corporations (Erramilli 1996). In terms of the cost, some pieces of additional information may require additional cost, or certification if pertaining to sensitive attributes.
@@
We summarize why some theoretical work on multi-objective contextual bandits cannot be directly applied to our problem formulation. In their study,  Tekin and Turğay (2018) addressed a contextual multi-armed bandit problem with two objectives, where one objective dominates the other. Their aim was to maximize the total reward in the non-dominant objective while ensuring that the dominant objective’s total reward is also maximized. However, our specific case requires the minimization of the total expected cost of the support policy while ensuring that the expected accuracy of the decision-maker under the policy remains above a certain threshold. Therefore, the techniques presented by here cannot be directly applied to our scenario. Turgay, Oner, and Tekin (2018) investigated the multi-objective contextual bandit problem with similarity information. Their approach relies on the assumption that a Lipschitz condition holds for the set of feasible context-arm pairs concerning the expected rewards for all objectives and that the learning algorithm has knowledge of the corresponding distance function (see Assumption 1 of their paper).
Appendix C Additional Details on Problem Formulation
Cost-agnostic Setting.
The optimization problem in the standard setting, where the only objective relates to expected loss, can be written as follows:
@@ Appendix C Additional Details on Problem Formulation
The human decision-making process with various forms of support can be effectively modeled as a stochastic contextual bandit problem. In this model, the diverse forms of support represent the available arms, and ${\mathcal{X}}$ represents the context space. For our investigation, we leverage two simple and efficient techniques from the existing contextual bandit literature: LinUCB (Li et al. 2010) and KNN-UCB (Guan and Jiang 2018). It’s worth noting that while any contextual bandit algorithm could be employed, we chose these two simpler methods as they have demonstrated their effectiveness in empirically showcasing the utility of our online decision support policy learning framework.
We present specific implementations of the Modiste algorithm (see Algorithm 1) using LinUCB and KNN-UCB in Algorithms 2 and 3, respectively. Importantly, these algorithms naturally inherit the regret bounds associated with LinUCB and KNN-UCB, as discussed below.
@@
1:  Input: trade-off parameter $\lambda$; human decision-maker $h$; cost function $c:{\mathcal{A}}\to[0,1]$; UCB parameter $\alpha$
@@
$\pi_{t+1}(x)~{}\leftarrow~{}$ $\operatorname*{arg\,min}_{a\in\mathcal{A}}\lambda\cdot\widehat{r}_{a,t}(x;h)+(1-\lambda)\cdot c(a)-\alpha\cdot\sqrt{x^{\top}(\mathbf{A}_{a})^{-1}x}$ (Step 2)
@@
11:  Output: policy $\pi_{\lambda}^{\text{alg}}\leftarrow\pi_{T+1}$
@@
1:  Input: trade-off parameter $\lambda$; human decision-maker $h$; cost function $c:{\mathcal{A}}\to[0,1]$; warm-up steps $W$; number of neighbours $K$; exploration parameter $\gamma$
@@
$\pi_{\text{knn}}(x)~{}\leftarrow~{}$ $\operatorname*{arg\,min}_{a\in\mathcal{A}}\lambda\cdot\widehat{r}_{a,t}(x;h)+(1-\lambda)\cdot c(a)$
@@
11:  Output: policy $\pi_{\lambda}^{\text{alg}}\leftarrow\pi_{T+1}$
$\lambda$ Selection Strategies
There are many ways to select the $\lambda$ to use on unseen decision-makers using a population of decision-makers. We outline three such strategies below. Note we use B in the main paper.
  1. A.
    Most likely $\lambda$. For each simulator, we identify the set of $\{\lambda_{\text{sim-}j}\}$, which yield policies that meet $L_{h_{\text{sim-}j}}^{\text{opt}}+\epsilon$. We then take the $\lambda$ that occurs the most often across simulators. If no policy meets the threshold for $h_{\text{sim-}j}$ , we select the policy closest to $L_{h_{\text{sim-}j}}^{\text{opt}}$, which can be computed given $r_{\textsc{A}_{i}}$ for all $x$ for each simulated decision-maker.
  2. B.
    Most likely $\lambda$ with lowest cost. For each simulator, we identify the set of $\{\lambda_{\text{sim-}j}\}$, which yield policies that meet $L_{h_{\text{sim-}j}}^{\text{opt}}+\epsilon$, and select the $\lambda$ whose policy has the least expected cost. Then, we identify the most common value across simulators. While the set of values here is a subset of the selection strategy A, a cost-aware selection strategy may identify a parameter that leads to a lower cost for new decision-makers.
  3. C.
    Conservative $\lambda$. For each simulator, we identify $\lambda_{\text{sim-}j}$ which is the minimum parameter that meets $L_{h_{\text{sim-}j}}^{\text{opt}}+\epsilon$. We then take $\max_{j}\lambda_{\text{sim-}j}$ over all simulated decision-makers. In the absence of suitable population-level information, this strategy is similar to picking a conservative value of $\lambda$ that prioritizes performance.
Appendix D Additional Experimental Details
@@
Two tasks (Synthetic-$2A$ and CIFAR-$2A$) are used only in the computational experiments, so we describe them abstractly. Both have the same set-up, but have different underlying data generating distributions. Consider learning a decision support policy in a setting where there are two forms of support available: $\mathcal{A}=\{\textsc{A}_{1},\textsc{A}_{2}\}$, hence CIFAR-$2A$. Let $c(\textsc{A}_{1})=0$ and $c(\textsc{A}_{2})=0.5$. We simulate a decision-maker’s behavior under each $r_{\textsc{A}_{i}}$ across 3 classes for Synthetic/CIFAR. We instantiate the $m$ decision-makers in the population data using human simulations, where we randomly sample $r_{\textsc{A}_{i}}(x;h)$ from a distribution for each $h$ in the population.
@@
We consider questions from four topics of MMLU (elementary mathematics, high school biology, high school computer science, and US foreign policy). Topic names match those proposed in (Hendrycks et al. 2020). We select topics to cover span an array of disciplines across the sciences and humanities, as discussed in Appendix F. We use questions in the MMLU test set for each topic that are at most 150 characters long. We limit the length of questions to facilitate readability in our human user studies and wanted to maximize parity between our computational and human experiments by considering the same set of questions across both set-ups. This yields 264, 197, 98, and 47 questions for the elementary mathematics, high school biology, US foreign policy, and high school computer science topics, respectively. We emphasize that in CIFAR, our forms of support operate in label space – in contrast, with MMLU, support selection is in covariate (topic) space – further highlighting the flexibility of our adaptive support paradigm.
@@
We consider a 5-class subset of CIFAR (Bird, Cat, Deer, Frog, and Horse), as discussed in Appendix E [^5] We explored a 3 class variant for HSEs to directly match the computational experiments; however, we realized that participants were able to figure out that which classes were impoverished, raising the base rate of correctly categorizing such images. Such behavior again highlights the need to carefully consider real human behavior in adaptive decision support systems.. We instantiate two forms of support: 1) a simulated AI model which provides a prediction for the image class, and 2) a consensus response derived from real humans, derived from the approximately 50 human annotators from CIFAR-10H (Peterson et al. 2019; Battleday, Peterson, and Griffiths 2020), presented as a distribution over the 5 classes.
We treat the original CIFAR-10 test set labels as the “true” labels and only include images for which the original CIFAR-10 label matches the label deemed most likely from the CIFAR-10H annotators (discrepancy was a rare occurrence, only 1.1% of the 5000 images considered). We sample a pool of 300 such images, and construct three different batches of 100 images. Participants are assigned to one batch of 100 images. Images per batch are sampled in a class-balanced fashion (i.e., 20 images for each of the 5 categories). Images are shuffled for each participant. The same latent codes ($z_{t}$) used in the computational experiments are employed for the user studies. For CIFAR, we define the support costs as: $c(\textsc{Human Alone})=0$ and $c(\textsc{Model})=c(\textsc{Consensus})=0.5$.
@@
Participants respond to 60 multiple choice questions, which were balanced by topic (15 questions per topic). Participants are assigned to one of three possible batches of 60 such questions. Order is shuffled. Participants are informed of the topic associated with each question (e.g., that the question was about biology). We implement a 10 second delay between when the question is presented and when the participant is allowed to submit their response to encourage participants to try each problem in earnest. For MMLU, we let the support costs be $c(\textsc{Human Alone})=0$ and $c(\textsc{Model})=0.1$.
@@
Figure 8: We plot the expected accuracy and the negative of the expected cost for an individual sampled from the Synthetic-$2A$ dataset for $\lambda\in[0,1]$. The ideal policy would lie in the far right corner. The red region denotes policies that fall within $\epsilon$ of the best risk for the sampled individual. We observe that there exists a policy that lies in the red region for both Modiste-KNN and Modiste-LinUCB. We also vary the cost structure (i.e., the cost of $\textsc{A}_{1}=0$ and the cost of $\textsc{A}_{2}=0.25,0.5,0.75$ in the first, second, and third row respectively).
Appendix E Additional Computational Experiments
Figure 9: Top is CIFAR-$3A$ and Bottom is MMLU-$2A$. For each of the $N=10$ simulators, we instantiate using the population data, we compute the difference in expected loss between the optimal policy for that simulator and the learned policy for all values of $\lambda\in[0,1]$ in increments of 0.05, and compute the difference in expected cost in the same manner. For KNN, we observe overlap between many values of $\lambda$ that all come within $\epsilon$ of the Best Risk: this implies we can use lower values of $\lambda$ to potentially achieve the desired level of performance. For LinUCB, we find that large values of $\lambda$ are required to achieve a suitable level of performance.
Implications of selecting the trade-off parameter $\lambda$ (with varying cost structure).
We follow the strategies to identify $\lambda$ for our cost-aware experiments. In Figure 9, we visualized the expected excess loss and cost for each value of $\lambda$ that is used for our human-informed synthetic decision-makers. Further, as shown in Figure 8, we highlight how only a subset of $\lambda$ values falls within $\epsilon=0.05$ of the $L_{h}^{\text{opt}}$ of that individual (e.g., in Figure 8). We also find that Modiste-KNN has more monotonic and controllable behavior as $\lambda$ increases, while $\lambda$ seems to have less of an effect on Modiste-LinUCB. We note that the set of $\lambda$ values that fall within the red region is affected by the choice of cost structure, as shown in Figure 8. Thus, any changes to the cost structure would imply the need to redo the hyper-parameter tuning process.
[…]
@@ Appendix E Additional Computational Experiments
Algorithm $L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$
Synthetic-$2A$ $\textsc{A}_{1}$ Only $0.51\pm 0.05$ $0.0$ $0.50\pm 0.06$ $0.0$ $0.50\pm 0.06$ $0.0$
$\textsc{A}_{2}$ Only $0.10\pm 0.03$ $0.5$ $0.30\pm 0.05$ $0.5$ $0.50\pm 0.04$ $0.5$
Population $0.37\pm 0.04$ $0.13\pm 0.03$ $0.34\pm 0.05$ $0.13\pm 0.02$ $0.49\pm 0.06$ $0.12\pm 0.03$
Modiste-LinUCB $\mathbf{0.15\pm 0.05}$ $\mathbf{0.31\pm 0.03}$ $0.12\pm 0.04$ $0.32\pm 0.03$ $0.34\pm 0.11$ $0.28\pm 0.07$
Modiste-KNN $0.11\pm 0.04$ $0.36\pm 0.03$ $\mathbf{0.13\pm 0.04}$ $\mathbf{0.30\pm 0.03}$ $\mathbf{0.33\pm 0.06}$ $\mathbf{0.21\pm 0.03}$
Best Risk $0.1$ $0.33$ $0.1$ $0.33$ $0.3$ $0.17$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.1,0.1,0.1]$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.1,0.7,0.1]$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.7,0.7,0.1]$
Algorithm $L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$ $L_{h}(\pi)$ $c(\pi)$
CIFAR–$2A$ $\textsc{A}_{1}$ Only $0.50\pm 0.05$ $0.0$ $0.50\pm 0.05$ $0.0$ $0.47\pm 0.05$ $0.0$
$\textsc{A}_{2}$ Only $0.1\pm 0.03$ $0.5$ $0.31\pm 0.04$ $0.5$ $0.49\pm 0.05$ $0.5$
Population $0.46\pm 0.07$ $0.03\pm 0.04$ $0.43\pm 0.06$ $0.07\pm 0.03$ $0.50\pm 0.06$ $0.03\pm 0.02$
Modiste-LinUCB $\mathbf{0.15\pm 0.04}$ $\mathbf{0.31\pm 0.02}$ $\mathbf{0.14\pm 0.05}$ $\mathbf{0.32\pm 0.02}$ $0.34\pm 0.07$ $0.25\pm 0.07$
Modiste-KNN $0.14\pm 0.04$ $0.33\pm 0.06$ $\mathbf{0.14\pm 0.04}$ $\mathbf{0.32\pm 0.03}$ $\mathbf{0.33\pm 0.05}$ $\mathbf{0.18\pm 0.03}$
Best Risk $0.1$ $0.33$ $0.1$ $0.33$ $0.3$ $0.17$
@@
Algorithm $r_{h}(\pi)$ $c(\pi)$ $r_{h}(\pi)$ $c(\pi)$ $r_{h}(\pi)$ $c(\pi)$
MMLU-$2A$ H-Only ($\pi^{*}_{\lambda=0}$) $0.45\pm 0.04$ $0.0$ $0.33\pm 0.06$ $0.0$ $0.51\pm 0.04$ $0.0$
H-Model $0.14\pm 0.02$ $0.1$ $0.30\pm 0.05$ $0.1$ $0.45\pm 0.04$ $0.1$
Human Pop $0.45\pm 0.07$ $0.05\pm 0.01$ $0.22\pm 0.08$ $0.06\pm 0.01$ $0.41\pm 0.04$ $0.05\pm 0.01$
Modiste-LinUCB $0.15\pm 0.04$ $0.05\pm 0.06$ $0.19\pm 0.07$ $0.05\pm 0.02$ $0.39\pm 0.08$ $0.06\pm 0.01$
Modiste-KNN $\mathbf{0.14\pm 0.06}$ $\mathbf{0.04\pm 0.01}$ $\mathbf{0.13\pm 0.04}$ $\mathbf{0.03\pm 0.04}$ $\mathbf{0.33\pm 0.05}$ $\mathbf{0.05\pm 0.01}$
Best Risk ($\pi^{*}_{\lambda=1}$) $0.13$ $0.05$ $0.13$ $0.03$ $0.3$ $0.05$
Selected $\lambda$ Values Algorithm $r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$ $r_{\textsc{A}_{2}}=[0.1,0.1,0.1]$ $r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$ $r_{\textsc{A}_{2}}=[0.1,0.7,0.1]$ $r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$ $r_{\textsc{A}_{2}}=[0.7,0.7,0.1]$ Synthetic-$2A$ Modiste-LinUCB $0.4$ $0.6$ $0.6$ Modiste-KNN $0.9$ $0.6$ $0.6$ CIFAR-$2A$ Modiste-LinUCB $0.4$ $0.9$ $0.5$ Modiste-KNN $0.7$ $0.9$ $0.6$
Algorithm $r_{\textsc{A}_{1}}=[0.8,0.8,0.1,0.1]$ $r_{\textsc{A}_{2}}=[0.2,0.1,0.1,0.2]$ $r_{\textsc{A}_{1}}=[0.2,0.8,0.1,0.1]$ $r_{\textsc{A}_{2}}=[0.8,0.1,0.1,0.2]$ $r_{\textsc{A}_{1}}=[0.2,0.5,0.5,0.8]$ $r_{\textsc{A}_{2}}=[0.8,0.1,0.1,0.8]$ MMLU-$2A$ Modiste-LinUCB $1.0$ $0.9$ $1.0$ Modiste-KNN $0.3$ $0.2$ $0.3$
Table 5: The best value of $\lambda$ that is selected from doing a sweep over $\lambda\in[0,1]$ that corresponds to the human simulators in Table 2. The column ordering is the same as the main paper.
Algorithm
Best Alone
Second Best Alone
Highest Best Risk
Lowest Best Risk
MMLU-$2A$ Modiste-LinUCB $0.7$ $0.65$ $1.0$ $0.95$
Modiste-KNN $0.25$ $0.55$ $0.35$ $0.6$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.1,0.1,0.1]$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.1,0.7,0.1]$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.7,0.7,0.1]$
Algorithm $r_{h}(\pi)$ $c(\pi)$ $r_{h}(\pi)$ $c(\pi)$ $r_{h}(\pi)$ $c(\pi)$
Synthetic-$2A$ Modiste-LinUCB ($\alpha=0.1$) $0.14\pm 0.04$ $0.33\pm 0.02$ $0.19\pm 0.09$ $0.32\pm 0.02$ $0.34\pm 0.10$ $0.21\pm 0.07$
Modiste-LinUCB ($\alpha=10$) $0.13\pm 0.04$ $0.32\pm 0.03$ $0.12\pm 0.03$ $0.33\pm 0.02$ $0.30\pm 0.05$ $0.23\pm 0.07$
Modiste-KNN ($\gamma=0.01$) $0.12\pm 0.04$ $0.35\pm 0.06$ $0.19\pm 0.08$ $0.26\pm 0.07$ $0.33\pm 0.04$ $0.20\pm 0.03$
Modiste-KNN ($\gamma=0.2$) $0.16\pm 0.05$ $0.33\pm 0.04$ $0.21\pm 0.06$ $0.24\pm 0.03$ $0.34\pm 0.06$ $0.15\pm 0.02$
CIFAR-$2A$ Modiste-LinUCB ($\alpha=0.1$) $0.15\pm 0.08$ $0.30\pm 0.04$ $0.14\pm 0.05$ $0.32\pm 0.02$ $0.31\pm 0.05$ $0.26\pm 0.07$
Modiste-LinUCB ($\alpha=10$) $0.14\pm 0.04$ $0.32\pm 0.02$ $0.18\pm 0.04$ $0.31\pm 0.02$ $0.32\pm 0.05$ $0.26\pm 0.08$
Modiste-KNN ($\gamma=0.01$) $0.12\pm 0.03$ $0.33\pm 0.02$ $0.14\pm 0.06$ $0.31\pm 0.03$ $0.35\pm 0.06$ $0.14\pm 0.03$
Modiste-KNN ($\gamma=0.2$) $0.17\pm 0.05$ $0.32\pm 0.06$ $0.20\pm 0.09$ $0.31\pm 0.03$ $0.35\pm 0.06$ $0.16\pm 0.04$
$r_{\textsc{A}_{1}}=[0.8,0.8,0.1,0.1]$
$r_{\textsc{A}_{2}}=[0.2,0.1,0.1,0.2]$
$r_{\textsc{A}_{1}}=[0.2,0.8,0.1,0.1]$
$r_{\textsc{A}_{2}}=[0.8,0.1,0.1,0.2]$
$r_{\textsc{A}_{1}}=[0.2,0.5,0.5,0.8]$
$r_{\textsc{A}_{2}}=[0.8,0.1,0.1,0.8]$
Algorithm $r_{h}(\pi)$ $c(\pi)$ $r_{h}(\pi)$ $c(\pi)$ $r_{h}(\pi)$ $c(\pi)$
MMLU-$2A$ Modiste-LinUCB ($\alpha=0.1$) $0.26\pm 0.06$ $0.03\pm 0.01$ $0.22\pm 0.09$ $0.04\pm 0.01$ $0.47\pm 0.08$ $0.03\pm 0.02$
Modiste-LinUCB ($\alpha=10$) $0.15\pm 0.05$ $0.04\pm 0.01$ $0.18\pm 0.07$ $0.03\pm 0.01$ $0.40\pm 0.06$ $0.05\pm 0.01$
Modiste-KNN ($\gamma=0.01$) $0.16\pm 0.05$ $0.05\pm 0.00$ $0.15\pm 0.06$ $0.03\pm 0.01$ $0.39\pm 0.06$ $0.03\pm 0.01$
Modiste-KNN ($\gamma=0.2$) $0.15\pm 0.03$ $0.05\pm 0.01$ $0.12\pm 0.03$ $0.03\pm 0.00$ $0.38\pm 0.05$ $0.03\pm 0.01$
Additional Experiments on Datasets.
In Table 4, we provide additional computational results for CIFAR-$2A$ and MMLU-$2A$ in the cost-aware setting. We find that Modiste successfully personalizes decision support policies that trade-off cost and performance effectively. We note that in all of our experiments, we fix the exploration parameters in LinUCB and KNN (i.e., $\alpha=1$ and $\gamma=0.1$ respectively). However, we do explore varying these exploration parameters in Table 6.
@@
Table 7: We vary the $K$ and $W$ parameters used to instantiate KNN across three datasets and report the expected risk $L_{h}(\pi)$ (lower is better) and expected cost $c(\pi)$ (lower is better) across 5 runs. The human simulation used here is the same as the one used in the second setting of Table 4.

### v3

@@ Appendix A Comparison Against Prior Work
  • CONSENSUS: One can also depict forms of support that are independent of any model, for instance, presenting the belief of one or more humans. Belief distributions can be constructed by pooling over many different humans’ “votes” for what a label ought to be, e.g., (Peterson et al. 2019; Beyer et al. 2020; Uma, Almanea, and Poesio 2022; Uma et al. 2020; Gordon et al. 2021, 2022), or by eliciting distributions over the likely label directly from each individual human (Collins, Bhatt, and Weller 2022; Collins et al. 2023). These consensus distributions permit the expression of uncertainty without any model. However, the elicitation of this form of support may be costly and humans may be fallible in the information they provide, e.g., due to direct labeling errors (Dawid and Skene 1979; Augustin et al. 2017; Whitehill et al. 2009; Wei et al. 2022), or miscalibrated confidence (O’Hagan et al. 2006; Collins et al. 2023; Lichtenstein, Fischhoff, and Phillips 1977; Tversky and Kahneman 1996).
  • ADDITIONAL: While much of this paper focused on support that provides decision-makers with label information, decision support may also entail acquiring or displaying additional contextual information (e.g., new features (Bakker et al. 2021)) or requesting previously unseen features, for instance, through additional medical diagnostics  (Harrell et al. 1982; Mylonakis et al. 2000). This flavor of support can be varied structurally, ranging from the results of a search query (Nakano et al. 2021) to hierarchical information like exposing the subsidiary ownership structure for multinational corporations (Erramilli 1996). Some pieces of additional information may require additional cost, or certification if pertaining to sensitive attributes.
@@
We summarize why some theoretical work on multi-objective contextual bandits cannot be directly applied to our problem formulation. In their study,  Tekin and Turğay (2018) addressed a contextual multi-armed bandit problem with two objectives, where one objective dominates the other. Their aim was to maximize the total reward in the non-dominant objective while ensuring that the dominant objective’s total reward is also maximized. However, our specific case requires the minimization of the total expected loss of the support policy while ensuring that the expected accuracy of the decision-maker under the policy remains above a certain threshold. Therefore, the techniques presented by here cannot be directly applied to our scenario. Turgay, Oner, and Tekin (2018) investigated the multi-objective contextual bandit problem with similarity information. Their approach relies on the assumption that a Lipschitz condition holds for the set of feasible context-arm pairs concerning the expected rewards for all objectives and that the learning algorithm has knowledge of the corresponding distance function (see Assumption 1 of their paper).
Appendix B Additional Details on Problem Formulation
The optimization problem in the standard setting, where the only objective relates to expected loss, can be written as:
@@ Appendix B Additional Details on Problem Formulation
The human decision-making process with various forms of support can be effectively modeled as a stochastic contextual bandit problem. In this model, the diverse forms of support represent the available arms, and ${\mathcal{X}}$ represents the context space. For our investigation, we leverage two simple and efficient techniques from the existing contextual bandit literature: LinUCB (Li et al. 2010) and KNN-UCB (Guan and Jiang 2018). It’s worth noting that while any contextual bandit algorithm could be employed, we chose these two simpler methods as they have demonstrated their effectiveness in empirically showcasing the utility of our online decision support policy learning framework. We present specific implementations of the Modiste algorithm (see Algorithm 1) using LinUCB and KNN-UCB in Algorithms 2 and 3, respectively.
@@
1:  Input: human decision-maker $h$; UCB parameter $\alpha$
@@
$\pi_{t+1}(x)~{}\leftarrow~{}$ $\operatorname*{arg\,min}_{a\in\mathcal{A}}\widehat{r}_{a,t}(x;h)-\alpha\cdot\sqrt{x^{\top}(\mathbf{A}_{a})^{-1}x}$ (Step 2)
@@
11:  Output: policy $\pi^{\text{alg}}\leftarrow\pi_{T+1}$
@@
1:  Input: human decision-maker $h$; warm-up steps $W$; number of neighbours $K$; exploration parameter $\gamma$
@@
$\pi_{\text{knn}}(x)~{}\leftarrow~{}$ $\operatorname*{arg\,min}_{a\in\mathcal{A}}\widehat{r}_{a,t}(x;h)$
@@
11:  Output: policy $\pi^{\text{alg}}\leftarrow\pi_{T+1}$
Appendix C Additional Experimental Details
@@
Two tasks (Synthetic-$2A$ and CIFAR-$2A$) are used only in the computational experiments, so we describe them abstractly. Both have the same set-up, but have different underlying data generating distributions. Consider learning a decision support policy in a setting where there are two forms of support available: $\mathcal{A}=\{\textsc{A}_{1},\textsc{A}_{2}\}$, hence CIFAR-$2A$. We simulate a decision-maker’s behavior under each $r_{\textsc{A}_{i}}$ across 3 classes for Synthetic/CIFAR. We instantiate the $m$ decision-makers in the population data using human simulations, where we randomly sample $r_{\textsc{A}_{i}}(x;h)$ from a distribution for each $h$ in the population.
@@
We consider questions from four topics of MMLU (elementary mathematics, high school biology, high school computer science, and US foreign policy). Topic names match those proposed in (Hendrycks et al. 2020). We select topics to cover span an array of disciplines across the sciences and humanities, as discussed in Appendix E. We use questions in the MMLU test set for each topic that are at most 150 characters long. We limit the length of questions to facilitate readability in our human user studies and wanted to maximize parity between our computational and human experiments by considering the same set of questions across both set-ups. This yields 264, 197, 98, and 47 questions for the elementary mathematics, high school biology, US foreign policy, and high school computer science topics, respectively. We emphasize that in CIFAR, our forms of support operate in label space – in contrast, with MMLU, support selection is in covariate (topic) space – further highlighting the flexibility of our adaptive support paradigm.
@@
We consider a 5-class subset of CIFAR (Bird, Cat, Deer, Frog, and Horse), as discussed in Appendix D [^4] We explored a 3 class variant for HSEs to directly match the computational experiments; however, we realized that participants were able to figure out that which classes were impoverished, raising the base rate of correctly categorizing such images. Such behavior again highlights the need to carefully consider real human behavior in adaptive decision support systems.. We instantiate two forms of support: 1) a simulated AI model which provides a prediction for the image class, and 2) a consensus response derived from real humans, derived from the approximately 50 human annotators from CIFAR-10H (Peterson et al. 2019; Battleday, Peterson, and Griffiths 2020), presented as a distribution over the 5 classes.
We treat the original CIFAR-10 test set labels as the “true” labels and only include images for which the original CIFAR-10 label matches the label deemed most likely from the CIFAR-10H annotators (discrepancy was a rare occurrence, only 1.1% of the 5000 images considered). We sample a pool of 300 such images, and construct three different batches of 100 images. Participants are assigned to one batch of 100 images. Images per batch are sampled in a class-balanced fashion (i.e., 20 images for each of the 5 categories). Images are shuffled for each participant. The same latent codes ($z_{t}$) used in the computational experiments are employed for the user studies.
@@
Participants respond to 60 multiple choice questions, which were balanced by topic (15 questions per topic). Participants are assigned to one of three possible batches of 60 such questions. Order is shuffled. Participants are informed of the topic associated with each question (e.g., that the question was about biology). We implement a 10 second delay between when the question is presented and when the participant is allowed to submit their response to encourage participants to try each problem in earnest.
@@
Appendix D Additional Computational Experiments
Figure 8: Snapshots of the learned decision support policies computed at the end of $100$ interactions with randomly sampled users for CIFAR-$3A$. The forms of support are colored in t-SNE embedding space. Both LinUCB and KNN generally identify the optimal form of support for a given input. All participant plots are included in Figure 12.
Algorithm
@@ Appendix D Additional Computational Experiments
Synthetic-$2A$ Modiste-LinUCB $\alpha=0.1$ $0.14\pm 0.04$ $0.19\pm 0.09$ $0.34\pm 0.10$
Modiste-LinUCB $\alpha=10$ $0.13\pm 0.04$ $0.12\pm 0.03$ $0.30\pm 0.05$
Modiste-KNN $\gamma=0.01$ $0.12\pm 0.04$ $0.19\pm 0.08$ $0.33\pm 0.04$
Modiste-KNN $\gamma=0.2$ $0.16\pm 0.05$ $0.21\pm 0.06$ $0.34\pm 0.06$
CIFAR-$2A$ Modiste-LinUCB $\alpha=0.1$ $0.15\pm 0.08$ $0.14\pm 0.05$ $0.31\pm 0.05$
Modiste-LinUCB $\alpha=10$ $0.14\pm 0.04$ $0.18\pm 0.04$ $0.32\pm 0.05$
Modiste-KNN $\gamma=0.01$ $0.12\pm 0.03$ $0.14\pm 0.06$ $0.35\pm 0.06$
Modiste-KNN $\gamma=0.2$ $0.17\pm 0.05$ $0.20\pm 0.09$ $0.35\pm 0.06$
Algorithm
@@
MMLU-$2A$ Modiste-LinUCB $\alpha=0.1$ $0.26\pm 0.06$ $0.22\pm 0.09$ $0.47\pm 0.08$
Modiste-LinUCB $\alpha=10$ $0.15\pm 0.05$ $0.18\pm 0.07$ $0.40\pm 0.06$
Modiste-KNN $\gamma=0.01$ $0.16\pm 0.05$ $0.15\pm 0.06$ $0.39\pm 0.06$
Modiste-KNN $\gamma=0.2$ $0.15\pm 0.03$ $0.12\pm 0.03$ $0.38\pm 0.05$
@@
Table 3: We vary the $K$ and $W$ parameters used to instantiate KNN across three datasets and report the expected risk $L_{h}(\pi)$ (lower is better) across 5 runs.